Circular No. 8761

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

RS OPHIUCHI M. F. Bode and M. J. Darnley, Liverpool John Moores University; J. P.

Osborne, A. P. Beardmore, M. R. Goad, and K. L. Page, University of Leicester; T. J. O’Brien and R. J. Davis, University of Manchester; A. Evans, Keele University; S. P. S. Eyres, University of Central Lancashire; S. Starrfield and J.-U. Ness, Arizona State University; G. Schwarz, West Chester University; J. Drake, Smithsonian Astrophysical Observatory; G. K. Skinner, Centre d’Etudes Spatiale des Rayonnements, Toulouse; N. Gehrels, Goddard Space Flight Center; and J. Krautter, University of Heidelberg, report that observations of the latest outburst of the recurrent nova RS Oph with the 2-m Liverpool Robotic Telescope on La Palma, Canary Islands, indicate a brightening that might be associated with the re-establishment of large-scale accretion from the red giant to the white dwarf in this system. Representative photometry of more fully sampled data is as follows: Sept. 8.92 UT (day 208 from outburst; cf. IAUC 8671), B = 13.61 ± 0.04, V = 12.16 ± 0.02, r′ = 10.83 ± 0.01, i′ = 10.06 ± 0.02, z′ = 9.50 ± 0.02; Sept. 19.89 (minor secondary peak in V, r′, i′, and z′; day 219), B = 13.19 ± 0.03, V = 11.43 ± 0.02, r′ = 10.19 ± 0.02, i′ = 9.53 ± 0.01, z′ = 9.11 ± 0.02; Oct. 11.88 (day 241, latest data), B = 13.07 ± 0.01, V = 11.49 ± 0.01, r′ = 10.27 ± 0.01, i′ = 9.60 ± 0.02, z′ = 9.13 ± 0.02. In addition, Swift continues to monitor the source (IAUC 8675, 8677; http://xxx.lanl.gov/abs/astro-

ph/?0604618; http://www.astronomerstelegram.org/?read=838, and

references therein). In observations taken roughly once every two weeks from days 112 to 190, the XRT 0.3- to 10-keV count rate decayed as a power law with index −3. However, observations on days 201 and 217 showed no flux decline, while the two most recent observations, on days 238 and 244, showed a decline with the same index. The XRT count rate at the flux standstill was 0.06 counts/s, corresponding to a luminosity of around 1.4 × 10^{33} erg/s. Both Swift and Liverpool Telescope observations are continuing. Complementary observations, particularly fast optical photometry and spectroscopy in the blue, to explore flickering of the central source, are encouraged.

COMET C/2006 M4 (SWAN) Visual total-magnitude estimates: Sept. 30.19 UT, 5.7 (N. Biver, Ver-

sailles, France, 7×50 binoculars); Oct. 6.21, 5.5 (J. Carvajal, Madrid, Spain, 6×30 binoculars); 14.81, 5.8 (J. J. Gonzalez, Leon, Spain, naked eye); 16.76, 6.0 (M. Meyer, Limburg, Germany, 10×50 binoculars).

2006 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green